OpenStudy (anonymous):

Write a rule for the linear function whose graph has slope m and contains the given point. m = 3; (-2, -2) a. y = 3x + 4 c. y = 3x + 2 b. y = 3x - 4 d. y = 3x - 2

OpenStudy (anonymous):

y-y(1)=m(x-x(1)) y-(-2)=3(x-(-2)) y+2=3(x+2) y+2=3x+6 -2 -2 y=3x+4

OpenStudy (anonymous):

thanks you ! :)